The length of a rectangle is 10ft longer than twice the width. If the perimeter is 92ft, find the length and width of the rectangle.
Answer by macston(5194)

W=width; L=length=2W+10ft; P=perimeter=2(L+W)=92 ft
P=2(L+W)
92 ft=2(L+W) Divide each side by 2.
46 ft=L+W Substitute for L.
46 ft=(2W+10 ft)+W Subtact 10 ft from each side.
36 ft=3W Divide each side by 3.
12 ft=W ANSWER: The width is 12 feet.
L=2W+10ft=2(12ft)+10ft=34ft ANSWER 2: The length is 34 feet.
CHECK:
P=2(L+W)
92 ft=2(34 ft+12 ft)
92 ft=2(46 ft)
92 ft=92 ft